# Sequence Reconstruction for Single-Deletion Single-Substitution Channel

### v1

@@ III-C For $\Lambda^{L}_{2}(\bm{x},\bm{x}^{\prime})$ and $\Lambda^{R}_{2}(\bm{x},\bm{x}^{\prime})$
For each $i\in\{1,2,\cdots,6\}$, let $\Lambda^{L}_{2,i}=\Lambda^{L}_{2,i}(\bm{x},\bm{x}^{\prime})$ be the set of $(x_{[n]\backslash j},x^{\prime}_{[n]\backslash j^{\prime}})\in\Lambda^{L}_{2}(\bm{x},\bm{x}^{\prime})$ such that $(j,j^{\prime})\in[n]\times[n]$, $j\leq j^{\prime}$ and conditions of Case $i$ hold. Then
$\Lambda^{L}_{2}=\bigcup_{i=1}^{6}\Lambda^{L}_{2,i}.$
If $|T^{L}\cap[1,i_{1}]|\geq 2$, let
@@
Similarly, the set $\Lambda^{R}_{2}=\Lambda^{R}_{2}(\bm{x},\bm{x}^{\prime})$ can be divided into six subsets $\Lambda^{R}_{2,i}=\Lambda^{R}_{2,i}(\bm{x},\bm{x}^{\prime})$, $i=1,2,\cdots,6~{}($see table 1)))), and
$\Lambda^{R}_{2}=\bigcup_{i=1}^{6}\Lambda^{R}_{2,i}.$
If $|T^{L}\cap[1,i_{1}]|\geq 2$, let
@@
Then
$\Omega^{X}_{2}=\bigcup_{i=1}^{6}\Omega^{X}_{2,i}.$
Moreover, we can easily obtain $\Omega^{X}_{2,i}$ from $\Lambda^{X}_{2,i}$ by Remark 2. As an example, consider $\Omega^{L}_{2,4}$ with the assumption of $|T^{L}\cap[i_{1}+1,i_{d-1}]|=0$. By Claim 2.4, we have $\Lambda^{L}_{2,4}\subseteq\{(x_{[n]\backslash k_{1}-1},x^{\prime}_{[n]\backslash i_{d-1}}),(x_{[n]\backslash i_{1}},x^{\prime}_{[n]\backslash j_{1}^{\prime}})\}$, where $j_{1}^{\prime}=\min(T^{L}\cap[i_{d-1}+1,i_{d}-1])$ when $T^{L}\cap[i_{d-1}+1,i_{d}-1]\neq\emptyset$. Then $\Omega^{L}_{2,4}=B^{\text{S}}_{1}(x_{[n]\backslash k_{1}-1},x^{\prime}_{[n]\backslash i_{d-1}})\cup B^{\text{S}}_{1}(x_{[n]\backslash i_{1}},x^{\prime}_{[n]\backslash j_{1}^{\prime}})$ and by Remark 2, we have $B^{\text{S}}_{1}(x_{[n]\backslash k_{1}-1},x^{\prime}_{[n]\backslash i_{d-1}})=\{\phi^{k_{1}-1}_{k_{1};x^{\prime}_{k_{1}-1}}(\bm{x}),\phi^{k_{1}-1}_{i_{d};x^{\prime}_{i_{d}}}(\bm{x})\}=\{\phi^{i_{d-1}}_{k_{1}-1;x_{k_{1}}}(\bm{x}^{\prime}),\phi^{i_{d-1}}_{i_{d};x_{i_{d}}}(\bm{x}^{\prime})\}$ and $B^{\text{S}}_{1}(x_{[n]\backslash i_{1}},x^{\prime}_{[n]\backslash j_{1}^{\prime}})=\{\phi^{i_{1}}_{i_{1};x^{\prime}_{j_{1}^{\prime}-1}}(\bm{x}),\phi^{i_{1}}_{i_{d};x^{\prime}_{i_{d}}}(\bm{x})\}=\{\phi^{j_{1}^{\prime}}_{j_{1}^{\prime}-1;x_{j_{1}^{\prime}}}(\bm{x}^{\prime}),\phi^{j_{1}^{\prime}}_{i_{d};x_{i_{d}}}(\bm{x}^{\prime})\}$. In particular, $|B^{\text{S}}_{1}(\bm{z},\bm{z}^{\prime})|\leq 2$ for each $(\bm{z},\bm{z}^{\prime})\in\Lambda^{X}_{2}$ and each $X\in\{L,R\}$.
@@
Let $k_{a}=\max(T^{L}\cap[1,i_{1}]\backslash\{k_{1},k_{2}\})$ if $|T^{L}\cap[1,i_{1}]|\geq 3$, and $k_{a}=1$ otherwise, where $k_{1}$ and $k_{2}$ are defined by (7) and (14), respectively. Let $k^{\prime}_{a}=\min(T^{L}\cap[i_{d}+1,n]\backslash\{k^{\prime}_{1},k^{\prime}_{2}\})-1$ if $|T^{L}\cap[i_{d}+1,n]|\geq 3$, and $k_{a}^{\prime}=n$ otherwise, where $k_{1}^{\prime}$ and $k^{\prime}_{2}$ are defined by (8) and (15), respectively. Then we have
$k_{a}<k_{2}<k_{1}\leq i_{1}<i_{d}<k^{\prime}_{1}<k^{\prime}_{2}\leq k^{\prime}_{a}.$ (18)
@@
Conversely, suppose $|T^{L}\cap[i_{1}+1,i_{d}]|\leq 2$. We have the following three cases (Cases a--c).
Case a: $|T^{L}\cap[i_{1}+1,i_{d}]|=2$. By (18), it is easy to see that $(j,j^{\prime})$ satisfying the conditions of Case 2 if and only if $k_{1}\leq j\leq i_{1}<i_{d}\leq j^{\prime}<k^{\prime}_{1}$. Similar to Claim 1.1, we can prove ((((by Lemma 1)))) that $x_{[k_{1},i_{1}]}~{}($resp. $x^{\prime}_{[i_{d},k^{\prime}_{1}-1]})$ is contained in a run of $\bm{x}~{}($resp. $\bm{x}^{\prime})$, which implies that $\Lambda^{L}_{2,2}=\{(x_{[n]\backslash j},x^{\prime}_{[n]\backslash j^{\prime}}):k_{1}\leq j\leq i_{1}<i_{d}\leq j^{\prime}<k^{\prime}_{1}\}=\{(x_{[n]\backslash i_{1}},x^{\prime}_{[n]\backslash i_{d}})\}$.
Case b: $|T^{L}\cap[i_{1}+1,i_{d}]|=1$. By (18), it is easy to see that $(j,j^{\prime})$ satisfying the conditions of Case 2 if and only if it satisfies one of the following two conditions: i) $k_{2}\leq j<k_{1}\leq i_{1}<i_{d}\leq j^{\prime}<k^{\prime}_{1}$; and ii) $k_{1}\leq j\leq i_{1}<i_{3}<k^{\prime}_{1}\leq j^{\prime}<k^{\prime}_{2}$. Hence, we have $\Lambda^{L}_{2,2}\subseteq\{(x_{[n]\backslash j},x^{\prime}_{[n]\backslash j^{\prime}}):\text{condition i) holds}\}\cup\{(x_{[n]\backslash j},x^{\prime}_{[n]\backslash j^{\prime}}):\text{condition ii) holds}\}=\{(x_{[n]\backslash k_{1}-1},x^{\prime}_{[n]\backslash i_{d}}),(x_{[n]\backslash i_{1}},x^{\prime}_{[n]\backslash k^{\prime}_{1}})\}$, where the equality is obtained by applying Lemma 1.
Case c: $|T^{L}\cap[i_{1}+1,i_{d}]|=0$. By (18), it is easy to see that $(j,j^{\prime})$ satisfying the conditions of Case 2 if and only if it satisfies one of the following three conditions: i) $k_{a}\leq j<k_{2}<k_{1}\leq i_{1}<i_{d}\leq j^{\prime}<k^{\prime}_{1}$; ii) $k_{2}\leq j<k_{1}\leq i_{1}<i_{d}<k^{\prime}_{1}\leq j^{\prime}<k^{\prime}_{2}$; and iii) $k_{1}\leq j\leq i_{1}<i_{d}<k_{1}^{\prime}<k^{\prime}_{2}\leq j^{\prime}\leq k^{\prime}_{a}$. Hence, we have $\Lambda^{L}_{2,2}\subseteq\big{\{}(x_{[n]\backslash j},x^{\prime}_{[n]\backslash j^{\prime}}):\text{condition i) holds}\big{\}}\cup\big{\{}(x_{[n]\backslash j},x^{\prime}_{[n]\backslash j^{\prime}}):\text{condition ii) holds}\big{\}}\cup\big{\{}(x_{[n]\backslash j},x^{\prime}_{[n]\backslash j^{\prime}}):\text{condition iii) holds}\big{\}}=\big{\{}(x_{[n]\backslash k_{2}-1},x^{\prime}_{[n]\backslash i_{d}})$, $(x_{[n]\backslash k_{1}-1},x^{\prime}_{[n]\backslash k^{\prime}_{1}})$, $(x_{[n]\backslash i_{1}},x^{\prime}_{[n]\backslash k^{\prime}_{2}})\big{\}}$, where the equality is obtained by applying Lemma 1.
@@
Suppose $|T^{L}\cap[i_{1}+1,i_{d-1}]|=0$. We are to prove $\Lambda^{L}_{2,4}\subseteq\{(x_{[n]\backslash k_{1}-1},x^{\prime}_{[n]\backslash i_{d-1}}),(x_{[n]\backslash i_{1}},x^{\prime}_{[n]\backslash j_{1}^{\prime}})\}$, where $j_{1}^{\prime}=\min(T^{L}\cap[i_{d-1}+1,i_{d}-1])$ when $T^{L}\cap[i_{d-1}+1,i_{d}-1]\neq\emptyset$. We need to consider the following Cases a and b.
Case a: $T^{L}\cap[i_{d-1}+1,i_{d}-1]\neq\emptyset$. Let $j_{2}^{\prime}=\min(T^{L}\cap[i_{d-1}+1,i_{d}]\backslash j_{1}^{\prime})$ if $T^{L}\cap[i_{d-1}+1,i_{d}]\neq\emptyset$, and $j_{2}^{\prime}=i_{d}$ otherwise. Then $(j,j^{\prime})$ satisfies the conditions of Case 4 if and only if one of the following two conditions holds: i) $k_{2}\leq j<k_{1}\leq i_{1}\leq i_{d-1}\leq j^{\prime}<j_{1}^{\prime}$; ii) $k_{1}\leq j\leq i_{1}\leq i_{d-1}<j_{1}^{\prime}\leq j^{\prime}<j_{2}^{\prime}$. Hence, we have $\Lambda^{L}_{2,4}\subseteq\{(x_{[n]\backslash j},x^{\prime}_{[n]\backslash j^{\prime}}):\text{condition i) holds}\}\cup\{(x_{[n]\backslash j},x^{\prime}_{[n]\backslash j^{\prime}}):\text{condition ii) holds}\}=\{(x_{[n]\backslash k_{1}-1},x^{\prime}_{[n]\backslash i_{d-1}}),(x_{[n]\backslash i_{1}},x^{\prime}_{[n]\backslash j^{\prime}_{1}})\}$, where the equality is obtained by applying Lemma 1.
Case b: $T^{L}\cap[i_{d-1}+1,i_{d}-1]=\emptyset$. It is easy to see that $(j,j^{\prime})$ satisfies the conditions of Case 4 if and only if $k_{2}\leq j<k_{1}\leq i_{1}\leq i_{d-1}\leq j^{\prime}<j_{1}^{\prime}$. Hence, we have $\Lambda^{L}_{2,4}=\{(x_{[n]\backslash j},x^{\prime}_{[n]\backslash j^{\prime}}):k_{2}\leq j<k_{1}\leq i_{1}\leq i_{d-1}\leq j^{\prime}<j_{1}^{\prime}\}=\{(x_{[n]\backslash k_{1}-1},x^{\prime}_{[n]\backslash i_{d-1}})\}$.
@@
Suppose $|T^{L}\cap[i_{2}+1,i_{d}]|=0$. We are to prove $\Lambda^{L}_{2,6}\subseteq\{(x_{[n]\backslash i_{2}},x^{\prime}_{[n]\backslash k^{\prime}_{1}}),(x_{[n]\backslash j_{1}-1},x^{\prime}_{[n]\backslash i_{d}})\}$, where $j_{1}=\max(T^{L}\cap[i_{1}+2,i_{2}])$ when $T^{L}\cap[i_{1}+2,i_{2}]\neq\emptyset$. We need to consider the following Cases a and b.
Case a: $T^{L}\cap[i_{1}+2,i_{2}]\neq\emptyset$. Let $j_{2}=\max(T^{L}\cap[i_{1}+2,i_{2}]\backslash j_{1})$ if $T^{L}\cap[i_{1}+2,i_{2}]\neq\emptyset$, and $j_{2}=i_{1}+1$ otherwise. Then $(j,j^{\prime})$ satisfies the conditions of Case 6 if and only if one of the following two conditions holds: i) $j_{1}\leq j\leq i_{2}\leq i_{d}<k_{1}^{\prime}\leq j^{\prime}<k_{2}^{\prime}$; ii) $j_{2}\leq j<j_{1}\leq i_{2}\leq i_{d}\leq j^{\prime}<k^{\prime}_{1}$. Hence, we have $\Lambda^{L}_{2,4}\subseteq\{(x_{[n]\backslash j},x^{\prime}_{[n]\backslash j^{\prime}}):\text{condition i) holds}\}\cup\{(x_{[n]\backslash j},x^{\prime}_{[n]\backslash j^{\prime}}):\text{condition ii) holds}\}=\{(x_{[n]\backslash i_{2}},x^{\prime}_{[n]\backslash k^{\prime}_{1}}),(x_{[n]\backslash j_{1}-1},x^{\prime}_{[n]\backslash i_{d}})\}$, where the equality is obtained by applying Lemma 1.
Case b: $T^{L}\cap[i_{1}+2,i_{2}]=\emptyset$. It is easy to see that $(j,j^{\prime})$ satisfies the conditions of Case 6 if and only if $j_{1}\leq j\leq i_{2}\leq i_{d}<k_{1}^{\prime}\leq j^{\prime}<k_{2}^{\prime}$. Hence, we have $\Lambda^{L}_{2,4}=\{(x_{[n]\backslash j},x^{\prime}_{[n]\backslash j^{\prime}}):j_{1}\leq j\leq i_{2}\leq i_{d}<k_{1}^{\prime}\leq j^{\prime}<k_{2}^{\prime}\}=\{(x_{[n]\backslash i_{2}},x^{\prime}_{[n]\backslash k^{\prime}_{1}})\}$.

### v2

@@ III-C For $\Lambda^{L}_{2}(\bm{x},\bm{x}^{\prime})$ and $\Lambda^{R}_{2}(\bm{x},\bm{x}^{\prime})$
For each $i\in\{1,2,\cdots,6\}$, let $\Lambda^{L}_{2,i}=\Lambda^{L}_{2,i}(\bm{x},\bm{x}^{\prime})$ be the set of $(x_{[n]\backslash j},x^{\prime}_{[n]\backslash j^{\prime}})\in\Lambda^{L}_{2}(\bm{x},\bm{x}^{\prime})$ such that $(j,j^{\prime})\in[n]\times[n]$, $j\leq j^{\prime}$ and conditions of Case $i$ hold. Then $\Lambda^{L}_{2}=\bigcup_{i=1}^{6}\Lambda^{L}_{2,i}.$ If $|T^{L}\cap[1,i_{1}]|\geq 2$, let
@@
Similarly, the set $\Lambda^{R}_{2}=\Lambda^{R}_{2}(\bm{x},\bm{x}^{\prime})$ can be divided into six subsets $\Lambda^{R}_{2,i}=\Lambda^{R}_{2,i}(\bm{x},\bm{x}^{\prime})$, $i=1,2,\cdots,6$, according to the value of $(|S\cap[1,j-1]|,|T^{R}\cap[j+1,j^{\prime}]|,|S\cap[j^{\prime}+1,n]|)~{}($see table 1)))), and $\Lambda^{R}_{2}=\bigcup_{i=1}^{6}\Lambda^{R}_{2,i}.$ If $|T^{L}\cap[1,i_{1}]|\geq 2$, let
@@
Then we have $\Omega^{X}_{2}=\bigcup_{i=1}^{6}\Omega^{X}_{2,i}.$ Moreover, we can easily obtain $\Omega^{X}_{2,i}$ from $\Lambda^{X}_{2,i}$ by Remark 2. As an example, consider $\Omega^{L}_{2,4}$ with the assumption of $|T^{L}\cap[i_{1}+1,i_{d-1}]|=0$. By Claim 2.4, we have $\Lambda^{L}_{2,4}\subseteq\{(x_{[n]\backslash k_{1}-1},x^{\prime}_{[n]\backslash i_{d-1}}),(x_{[n]\backslash i_{1}},x^{\prime}_{[n]\backslash j_{1}^{\prime}})\}$, where $j_{1}^{\prime}=\min(T^{L}\cap[i_{d-1}+1,i_{d}-1])$ when $T^{L}\cap[i_{d-1}+1,i_{d}-1]\neq\emptyset$. Then $\Omega^{L}_{2,4}=B^{\text{S}}_{1}(x_{[n]\backslash k_{1}-1},x^{\prime}_{[n]\backslash i_{d-1}})\cup B^{\text{S}}_{1}(x_{[n]\backslash i_{1}},x^{\prime}_{[n]\backslash j_{1}^{\prime}})$ and by Remark 2, we have $B^{\text{S}}_{1}(x_{[n]\backslash k_{1}-1},x^{\prime}_{[n]\backslash i_{d-1}})=\{\phi^{k_{1}-1}_{k_{1};x^{\prime}_{k_{1}-1}}(\bm{x}),\phi^{k_{1}-1}_{i_{d};x^{\prime}_{i_{d}}}(\bm{x})\}=\{\phi^{i_{d-1}}_{k_{1}-1;x_{k_{1}}}(\bm{x}^{\prime}),\phi^{i_{d-1}}_{i_{d};x_{i_{d}}}(\bm{x}^{\prime})\}$ and $B^{\text{S}}_{1}(x_{[n]\backslash i_{1}},x^{\prime}_{[n]\backslash j_{1}^{\prime}})=\{\phi^{i_{1}}_{i_{1};x^{\prime}_{j_{1}^{\prime}-1}}(\bm{x}),\phi^{i_{1}}_{i_{d};x^{\prime}_{i_{d}}}(\bm{x})\}=\{\phi^{j_{1}^{\prime}}_{j_{1}^{\prime}-1;x_{j_{1}^{\prime}}}(\bm{x}^{\prime}),\phi^{j_{1}^{\prime}}_{i_{d};x_{i_{d}}}(\bm{x}^{\prime})\}$. In particular, $|B^{\text{S}}_{1}(\bm{z},\bm{z}^{\prime})|=2$ for each $(\bm{z},\bm{z}^{\prime})\in\Lambda^{X}_{2}$ and each $X\in\{L,R\}$.
@@
The proofs of Claims 2.1--2.6 are similar to the proofs of Claims 1.1--1.3. Let
$k_{c}=\!\left\{\!\begin{aligned} &\max(T^{L}\cap[1,i_{1}]\backslash\{k_{1},k_{2}\}),~{}\text{if}~{}|T^{L}\cap[1,i_{1}]|\geq 3;\\ &1,~{}\text{otherwise}.\end{aligned}\right.$ (20)
and
$k^{\prime}_{c}=\!\left\{\!\begin{aligned} &\min(T^{L}\cap[i_{d}+1,n]\backslash\{k^{\prime}_{1},k^{\prime}_{2}\})-1,~{}\text{if}~{}|T^{L}\cap[i_{d}+1,n]|\geq 3;\\ &n,~{}\text{otherwise}.\end{aligned}\right.$ (21)
Here, $k_{1}$, $k_{1}^{\prime}$ $k_{2}$ and $k_{2}^{\prime}$ are defined by (7), (8), (16) and (17) respectively. Then we have
$k_{c}<k_{2}<k_{1}\leq i_{1}<i_{d}<k^{\prime}_{1}<k^{\prime}_{2}\leq k^{\prime}_{c}.$ (22)
@@
Conversely, suppose $|T^{L}\cap[i_{1}+1,i_{d}]|\leq 2$. We have the following Cases (i)--(iii).
Case (i): $|T^{L}\cap[i_{1}+1,i_{d}]|=2$. By (22), it is easy to see that $(j,j^{\prime})$ satisfying the conditions of Case 2 if and only if $k_{1}\leq j\leq i_{1}<i_{d}\leq j^{\prime}<k^{\prime}_{1}$. Similar to Claim 1.1, we can prove ((((by Lemma 1)))) that $x_{[k_{1},i_{1}]}~{}($resp. $x^{\prime}_{[i_{d},k^{\prime}_{1}-1]})$ is contained in a run of $\bm{x}~{}($resp. $\bm{x}^{\prime})$, which implies that $\Lambda^{L}_{2,2}=\{(x_{[n]\backslash j},x^{\prime}_{[n]\backslash j^{\prime}}):k_{1}\leq j\leq i_{1}<i_{d}\leq j^{\prime}<k^{\prime}_{1}\}=\{(x_{[n]\backslash i_{1}},x^{\prime}_{[n]\backslash i_{d}})\}$.
Case (ii): $|T^{L}\cap[i_{1}+1,i_{d}]|=1$. By (22), it is easy to see that $(j,j^{\prime})$ satisfying the conditions of Case 2 if and only if it satisfies one of the following two conditions: 1) $k_{2}\leq j<k_{1}\leq i_{1}<i_{d}\leq j^{\prime}<k^{\prime}_{1}$; and 2) $k_{1}\leq j\leq i_{1}<i_{3}<k^{\prime}_{1}\leq j^{\prime}<k^{\prime}_{2}$. Hence, we have $\Lambda^{L}_{2,2}\subseteq\{(x_{[n]\backslash j},x^{\prime}_{[n]\backslash j^{\prime}}):\text{condition i) holds}\}\cup\{(x_{[n]\backslash j},x^{\prime}_{[n]\backslash j^{\prime}}):\text{condition ii) holds}\}=\{(x_{[n]\backslash k_{1}-1},x^{\prime}_{[n]\backslash i_{d}}),(x_{[n]\backslash i_{1}},x^{\prime}_{[n]\backslash k^{\prime}_{1}})\}$, where the equality is obtained by applying Lemma 1.
Case (iii): $|T^{L}\cap[i_{1}+1,i_{d}]|=0$. By (22), it is easy to see that $(j,j^{\prime})$ satisfying the conditions of Case 2 if and only if it satisfies one of the following three conditions: 1) $k_{c}\leq j<k_{2}$ and $i_{d}\leq j^{\prime}<k^{\prime}_{1}$ (see Fig. 6 (a)), which implies $(x_{[n]\backslash k_{2}-1},x^{\prime}_{[n]\backslash i_{d}})\in\Lambda^{L}_{2,2}$; 2) $k_{2}\leq j<k_{1}$ and $k^{\prime}_{1}\leq j^{\prime}<k^{\prime}_{2}$ (see Fig. 6 (b)), which implies $(x_{[n]\backslash k_{1}-1},x^{\prime}_{[n]\backslash k^{\prime}_{1}})\in\Lambda^{L}_{2,2}$; and 3) $k_{1}\leq j\leq i_{1}$ and $k^{\prime}_{2}\leq j^{\prime}\leq k^{\prime}_{c}$ (see Fig. 6 (c)), which implies $(x_{[n]\backslash i_{1}},x^{\prime}_{[n]\backslash k^{\prime}_{2}})\in\Lambda^{L}_{2,2}$. Hence, we have $\Lambda^{L}_{2,2}\subseteq\big{\{}(x_{[n]\backslash k_{2}-1},x^{\prime}_{[n]\backslash i_{d}})$, $(x_{[n]\backslash k_{1}-1},x^{\prime}_{[n]\backslash k^{\prime}_{1}})$, $(x_{[n]\backslash i_{1}},x^{\prime}_{[n]\backslash k^{\prime}_{2}})\big{\}}$.
Refer to captionRefer to caption
Figure 6: An illustration of the pair $(j,j^{\prime})$ in Case (iii) of the proof of Claim 2.2. Here $S=\{i_{1},i_{2},i_{3}\}$. In this figure, (a) is for condition 1), (b) is for condition 2) and (c) is for condition 3).
@@
Suppose $|T^{L}\cap[i_{1}+1,i_{d-1}]|=0$. We are to prove $\Lambda^{L}_{2,4}\subseteq\{(x_{[n]\backslash k_{1}-1},x^{\prime}_{[n]\backslash i_{d-1}}),(x_{[n]\backslash i_{1}},x^{\prime}_{[n]\backslash j_{1}^{\prime}})\}$, where $j_{1}^{\prime}=\min(T^{L}\cap[i_{d-1}+1,i_{d}-1])$ when $T^{L}\cap[i_{d-1}+1,i_{d}-1]\neq\emptyset$. We need to consider the following Cases (i) and (ii).
Case (i): $T^{L}\cap[i_{d-1}+1,i_{d}-1]\neq\emptyset$. Let $j_{2}^{\prime}=\min(T^{L}\cap[i_{d-1}+1,i_{d}]\backslash j_{1}^{\prime})$ if $T^{L}\cap[i_{d-1}+1,i_{d}]\neq\emptyset$, and $j_{2}^{\prime}=i_{d}$ otherwise. Then $(j,j^{\prime})$ satisfies the conditions of Case 4 if and only if one of the following two conditions holds: 1) $k_{2}\leq j<k_{1}\leq i_{1}\leq i_{d-1}\leq j^{\prime}<j_{1}^{\prime}$; 2) $k_{1}\leq j\leq i_{1}\leq i_{d-1}<j_{1}^{\prime}\leq j^{\prime}<j_{2}^{\prime}$. Hence, we have $\Lambda^{L}_{2,4}\subseteq\{(x_{[n]\backslash j},x^{\prime}_{[n]\backslash j^{\prime}}):\text{condition 1) holds}\}\cup\{(x_{[n]\backslash j},x^{\prime}_{[n]\backslash j^{\prime}}):\text{condition 2) holds}\}=\{(x_{[n]\backslash k_{1}-1},x^{\prime}_{[n]\backslash i_{d-1}}),(x_{[n]\backslash i_{1}},x^{\prime}_{[n]\backslash j^{\prime}_{1}})\}$, where the equality is obtained by applying Lemma 1.
Case (ii): $T^{L}\cap[i_{d-1}+1,i_{d}-1]=\emptyset$. It is easy to see that $(j,j^{\prime})$ satisfies the conditions of Case 4 if and only if $k_{2}\leq j<k_{1}\leq i_{1}\leq i_{d-1}\leq j^{\prime}<j_{1}^{\prime}$. Hence, we have $\Lambda^{L}_{2,4}=\{(x_{[n]\backslash j},x^{\prime}_{[n]\backslash j^{\prime}}):k_{2}\leq j<k_{1}\leq i_{1}\leq i_{d-1}\leq j^{\prime}<j_{1}^{\prime}\}=\{(x_{[n]\backslash k_{1}-1},x^{\prime}_{[n]\backslash i_{d-1}})\}$.
@@
Suppose $|T^{L}\cap[i_{2}+1,i_{d}]|=0$. We are to prove $\Lambda^{L}_{2,6}\subseteq\{(x_{[n]\backslash i_{2}},x^{\prime}_{[n]\backslash k^{\prime}_{1}}),(x_{[n]\backslash j_{1}-1},x^{\prime}_{[n]\backslash i_{d}})\}$, where $j_{1}=\max(T^{L}\cap[i_{1}+2,i_{2}])$ when $T^{L}\cap[i_{1}+2,i_{2}]\neq\emptyset$. We need to consider the following Cases (i) and (ii).
Case (i): $T^{L}\cap[i_{1}+2,i_{2}]\neq\emptyset$. Let $j_{2}=\max(T^{L}\cap[i_{1}+2,i_{2}]\backslash j_{1})$ if $T^{L}\cap[i_{1}+2,i_{2}]\neq\emptyset$, and $j_{2}=i_{1}+1$ otherwise. Then $(j,j^{\prime})$ satisfies the conditions of Case 6 if and only if one of the following two conditions holds: 1) $j_{1}\leq j\leq i_{2}\leq i_{d}<k_{1}^{\prime}\leq j^{\prime}<k_{2}^{\prime}$; 2) $j_{2}\leq j<j_{1}\leq i_{2}\leq i_{d}\leq j^{\prime}<k^{\prime}_{1}$. Hence, we have $\Lambda^{L}_{2,4}\subseteq\{(x_{[n]\backslash j},x^{\prime}_{[n]\backslash j^{\prime}}):\text{condition 1) holds}\}\cup\{(x_{[n]\backslash j},x^{\prime}_{[n]\backslash j^{\prime}}):\text{condition 2) holds}\}=\{(x_{[n]\backslash i_{2}},x^{\prime}_{[n]\backslash k^{\prime}_{1}}),(x_{[n]\backslash j_{1}-1},x^{\prime}_{[n]\backslash i_{d}})\}$, where the equality is obtained by applying Lemma 1.
Case (ii): $T^{L}\cap[i_{1}+2,i_{2}]=\emptyset$. It is easy to see that $(j,j^{\prime})$ satisfies the conditions of Case 6 if and only if $j_{1}\leq j\leq i_{2}\leq i_{d}<k_{1}^{\prime}\leq j^{\prime}<k_{2}^{\prime}$. Hence, we have $\Lambda^{L}_{2,4}=\{(x_{[n]\backslash j},x^{\prime}_{[n]\backslash j^{\prime}}):j_{1}\leq j\leq i_{2}\leq i_{d}<k_{1}^{\prime}\leq j^{\prime}<k_{2}^{\prime}\}=\{(x_{[n]\backslash i_{2}},x^{\prime}_{[n]\backslash k^{\prime}_{1}})\}$.